Q.3.2 The store selling rice makes a further discovery, when the price of couscous changes from R30 per kg to R26 per kg, then the quantity of rice demanded decreases from 1360 kg per month to 1238 kg per month. Use this information to answer Q.3.2.1 and Q.3.2.2 below.

Q.3.2.1 Use the point method to calculate the cross elasticity of demand for these products.

Q.3.2.2 Comment on the significance of your solution in Q.3.2.1.

Q.3.3 Identify three factors that are likely to affect the price elasticity of supply for rice. Clearly indicate the impact that each factor will have on the price elasticity.


1
Expert's answer
2021-05-30T14:15:02-0400

Q.3.2.1

Mid Point Elasticity formula is given as below.

"E_d =\frac{ (Q_2-Q_1)}{(Q1+Q2)}\times \frac{(P1+P2)}{(P2 -P1)}"


"=\frac{ (1238-1360)}{(1238+1360)}\times \frac{(30+26)}{(26 -30)}"


"=\frac{-122}{2598}\times\frac{56}{-4}"


"=0.657428791"


Q.3.2.2

The elasticity of demand is positive meaning an increase in price increases the quantity demanded hence the prices should be set high. Therefore couscous is a veblen good.


Q.3.3

1. Availability of factors of productivity

When factors of production are available, supply will be highly elastic and vice versa. This is because when production factors are available the suppliers will be able to meet demand in good times.

2. Buffer-stock

If there exist a lot of stock, in case prices increase, supplier would be able to respond very fast by increasing supply. In such a case supply is said to be highly elastic.

3. Time factor

This refers to the time it takes to produce and supply a product in the market. In the short run, supply of rice will be inelastic. But, in the long run supply is elastic.